OpenStudy (anonymous):

Find the angle θ between the vectors. (Round your answer to two decimal places.) u = 20i + 40j v = −3j + 9k the answer i got is 31.95 but its wrong...

OpenStudy (anonymous):

31.95 degree

OpenStudy (mathmale):

What course are you in? What vector operation are you using to obtain the cosine of the angle between the two vectors?

OpenStudy (anonymous):

20 * -3 +40 * -3 then you take ||u|| * ||v|| then you simply use inverse of cos to get the answer

OpenStudy (mathmale):

Thank you for sharing your result (31.95 degrees), but I'm far more interested in how you obtained that.

OpenStudy (mathmale):

You're basically on the right track! But could you fix up the grammar and math involved in 20 * -3 +40 * -3 then you take ||u|| * ||v|| ? What do you mean by "you take ... " ?

OpenStudy (anonymous):

360/sqareroot of 2000 * 3 square root of 10 then just used the inverse of cos

OpenStudy (mathmale):

Parth, supposing you were seeing the dot product for the first time, and trying to apply it. Would you, in that situation, likely understand "20 * -3 +40 * -3 then you take ||u|| * ||v|| "??

OpenStudy (anonymous):

this is what i did first thing you have to find dot product of U dot V for dot product my result is 360. now i have to find ||U|| the answer for that is Sqareroot of 2000 or 44.72 now i have to find ||V|| the answer for that is 3 square root of 10 then i did 360/squareroot of 2000 times 3 square root of 10 answer for the above steps = 0.848528137 then inverse of cos of 0.848528137

OpenStudy (anonymous):

and thats how i got 31.95

OpenStudy (mathmale):

What I'm asking you to do, Parth, is to type out the exact formula that we need here. It is\[\cos \theta=\frac{ u~\dot~v }{ |u||v| }\] I have trouble finding DIVISION within your explanation.

OpenStudy (anonymous):

yes thats the one

OpenStudy (anonymous):

sorry i don't really know how to put formula like that

OpenStudy (anonymous):

i just joined today so

OpenStudy (mathmale):

Therefore, you'd have this result:\[\cos \theta=\frac{ u~\dot~v }{ |u||v| }=\frac{ <20,40><-3,9> }{ |u||v| }\]

OpenStudy (anonymous):

yes and i got 360 for the top part

OpenStudy (mathmale):

Welcome aboard OpenStudy!!! Have you already done the dot product indicated in the numerator? If so, what is that dot product?

OpenStudy (mathmale):

How did you get 360? My result is -60+360, or ... ?

OpenStudy (anonymous):

wait it should be 300 sorry

OpenStudy (mathmale):

Much better. So, now, Parth, you have the scalar quantity 300 in the numerator. What is the magnitude of vector u, please?

OpenStudy (anonymous):

the answer for the question should be 45 degree?

OpenStudy (anonymous):

square root of 2000

OpenStudy (anonymous):

and for V = 3 square root of 10

OpenStudy (mathmale):

Yes, and what is the magnitude of vector v?

OpenStudy (anonymous):

\[3\sqrt{10}\]

OpenStudy (anonymous):

45 degree is wrong answer

OpenStudy (mathmale):

Looks better now. Multiply those two magnitudes together, please. Let's not focus on the answer yet, ok, but rather focus on how we get there accurately.

OpenStudy (anonymous):

424.2640687

OpenStudy (anonymous):

thank you for taking your time to help!

OpenStudy (mathmale):

I'm going to take your word for that now, since it looks like it's in the correct ball park! So, the cosine of your angle is 300 over 424.3, right? What is the angle in degrees? In radians?

OpenStudy (mathmale):

Are you located in India? If so, in which city?

OpenStudy (anonymous):

degree

OpenStudy (anonymous):

no, i am in Texas, USA

OpenStudy (mathmale):

Parth, let's try those absolute values once more.

OpenStudy (mathmale):

You got a denom. of 424 something, whereas mine was 684 something. We have to be 100% in agreement on this. vector u is <40,60>, right? double check the original problem statement.

OpenStudy (anonymous):

u = 20i + 40j so it should be <20, 40>

OpenStudy (mathmale):

The square of 40 is 1600 and that of 60 is 3600, correct? Adding those squares together, we get 5200, correct?

OpenStudy (mathmale):

If so, just find the amplitude of u as a mixed decimal number.

OpenStudy (mathmale):

For vector v: the magnitude is the square root of 9+81, or square root of 90, or 3Sqrt(10). I'm in agreement with you there.

OpenStudy (anonymous):

yeah but for U i have <20, 40> so 20^2 + 40^2 = square root of 2000

OpenStudy (mathmale):

Yes, that looks correct, so the denominator is Sqrt(2000) time 3 times Sqrt (10). Does that boil down to about 424? If so, you're on the right track and I'm wrong.

OpenStudy (anonymous):

yes 424.26

OpenStudy (mathmale):

so the cosine of your angle is 300/424.26. What is your angle, in degrees?

OpenStudy (anonymous):

300/424.26 =0.707113562

OpenStudy (anonymous):

then inverse of cos (0.707113562) = 45 degree

OpenStudy (mathmale):

Right!! Great work. My advice: slow down a little; check everything twice.

OpenStudy (anonymous):

but the answer is wrong LMAO

OpenStudy (anonymous):

its online homework and i tried 45 and it gave me wrong answer

OpenStudy (mathmale):

What are your answer choices? Are they expressed in degrees or in radians?

OpenStudy (anonymous):

i don't have answer choice but its asking in degree

OpenStudy (mathmale):

(Thinking).

OpenStudy (anonymous):

just hold on me take a screenshot and post it here

OpenStudy (mathmale):

Hey, could you possibly take a screen shot of the original problem and share the image with me thru openstudy?

OpenStudy (mathmale):

Great minds think alike!!

OpenStudy (anonymous):

haha yeah there you go

OpenStudy (anonymous):

see my first answer was 45 but i guess its wrong

OpenStudy (mathmale):

I've done a quick, rough sketch of these two vectors; from that sketch it does seem that the angle is in the ballpark of 45 degrees. Let me look at your screenshot.

OpenStudy (anonymous):

Thank you so much

OpenStudy (mathmale):

Wow. As you know, "a picture is worth a thousand words."

OpenStudy (mathmale):

The reason why we're getting the wrong answer is that these vectors are in 3-d space, not in the plane. the first one is <20,40,0> and the second is <0,-3,9>. Are you able to find the dot product when you have vectors in 3-d space?

OpenStudy (anonymous):

uhh.. idk where did i go worng!

OpenStudy (mathmale):

Try finding the dot product <20,40,0> . <0,-3,9>.

OpenStudy (anonymous):

OHHHHHHHHHHHHH

OpenStudy (anonymous):

yeah its j, h, k thing

OpenStudy (mathmale):

We have the magnitudes of the vectors correct, so try letting cos theta =-120/424 something.

OpenStudy (anonymous):

i mean I J K

OpenStudy (mathmale):

Yes, a i, j, k thing, otherwise known as 3-d space! ;)

OpenStudy (anonymous):

the answer should be 106.44?

OpenStudy (anonymous):

LOL!! thank you so much, i didn't look at the alphabets hahah

OpenStudy (mathmale):

That's exactly what I got. Congrats!!!!

OpenStudy (mathmale):

You are hereby condemned to look at the alphabet all night.

OpenStudy (anonymous):

let me try that answer

OpenStudy (mathmale):

Always a good idea to check. Curious, though, just how you'll check that answer.

OpenStudy (anonymous):

its wrong and i am out of tries now so i have to talk with the professor to get it fixed again. thank you so much!!

OpenStudy (mathmale):

I think you've learned a lot by going thru this. I'd suggest you try drawing these 2 vectors in space and see whether that gives you a clue as to what the angle should be.

OpenStudy (anonymous):

yeah thank you!

OpenStudy (mathmale):

By all means please come back to OpenStudy. I look forward to working with you again! Good night!